15. Integrating Problem Samantha Roberts has a job as a pharmacist earning $30,000 per year, and she is deciding whether to take another job as the manager of another pharmacy for $40,000 per year or to purchase a pharmacy that generates a revenue of $200,000 per year. To purchase the pharmacy, Samantha would have to use her $20,000 savings and borrow another $80,000 at an interest rate of 10 percent per year. The pharmacy that Samantha is contemplating purchasing has additional expenses of $80,000 for supplies, $40,000 for hired help, $10,000 for rent, and $5,000 for utilities. Assume that income and business taxes are zero and that the repayment of the principal of the loan does not start before three years. (a) What would be the business and economic profit if Samantha purchased the pharmacy? Should Samantha purchase the pharmacy? (b) Suppose that Samantha expects that another pharmacy will open nearby at the end of three years and that this will drive the economic profit of the pharmacy to zero. What would the revenue of the pharmacy be in three years? (c) What theory of profit would account for profits being earned by the pharmacy during the first three years of operation? (d) Suppose that Samantha expects to sell the pharmacy at the end of three years for $50,000 more than the price she paid for it and that she requires a 15 percent return on her investment. Should she still purchase the pharmacy?
Note: 1. P15(d): Change ... for $50,000 more than ..to . for $50,000 less than ... Compare the present value of economic profit in each of the next three years and the loss of $50,000 in the third year using 15% as the discount rate.
Answer #1

Answer a)

Business profit = total revenue – total business cost (or explicit cost)
= revenue – supplies – interest on loan – utilities – hired help – rent
= 200000- 80000 - (80000*0.1) – 5000 – 40000 -10000
= $57000
Economic profit = business profit - interest rate forgone on saving amount – opportunity cost of purchasing pharmacy
= 57000 - (20000*0.1) - 40000= $15000
Samantha should purchase the pharmacy because economic profit is positive


Answer b)
If revenue is equal to economic cost economic profit would be zero.
Economic cost = interest on loan + supplies + rent + utilities + interest rate forgone on saving amount + hired help + opportunity cost of purchasing pharmacy = $185000
Revenue per year would be $185000


Answer c)
In the given case Samantha Jones is having no competitors in initial three years. She is operating alone and there is no nearby competitor. It is theory of monopoly in which Samantha earned handsome business and economic profits.


Answer d)
If Samantha sells the pharmacy after three years she is expected to get
50000+20000*0.15+80000*(Interest to be received - interest to be paid)
= 50000+20000*0.15+80000*(0.15-0.10) = $57000
If Samantha had worked as manager, she could have earned $120000 (40000*3) after 3 years.
Hence, Samantha should work as manager instead of purchasing the pharmacy.